Evaluate the function at the specified value of the independent variable and simplify.

f(x) = { x, x/= -1
x^2-3x, -1< x <= 1
x^3-3x^2, x>1
f(-2/5)

Answer by stanbon(75887)

Evaluate the function at the specified value of the independent variable and simplify.
f(x) = { x, x/= -1
f(-1) = -1
--------------------------
x^2-3x, -1< x <= 1
f(1) = 1-3 = -2
-------------------------
x^3-3x^2, x>1
f(x>1)> 1-3
f(x>1)> -2
-------------------------
f(-2/5)
Not sure what you are looking for here.
===========================================
Cheers,
Stan H.